OpenStudy (anonymous):

Help?(:

OpenStudy (misty1212):

lol HI!! (again)

OpenStudy (anonymous):

Katie wants to hang a painting in a gallery. The painting and frame must have an area of 45 square feet. The painting is 6 feet wide by 7 feet long. Which quadratic equation can be used to determine the thickness of the frame, x? (4 points) 4x2 + 26x + 45 = 0 4x2 + 26x - 3 = 0 x2 + 13x - 3 = 0 x2 + 13x + 45 = 0

OpenStudy (misty1212):

another quadratic equation?

OpenStudy (anonymous):

nope.

OpenStudy (misty1212):

looks like four actually

OpenStudy (anonymous):